You analyze the voltages across different branches of a simple circuit (using Kirchoffs Laws), and you find the following equations relating the voltages across different resistors 4V 1+4V 2 90 6.1 V 2+7 v 3 72 where vi represents current through wire 1 2 through wire 2, etc. All voltages are in V (Volts). Using matrix methods, determine the magnitude of vi1 in Volts Answer Check
Answer #1

The following equations relating the voltages across different resistors which are given below as :

# (4) V1 + (4) V2 = (90 V)

V1 = [(90 V) - (4) V2] / (4)                                               

V1 = [(45 V) - (2) V2] / (2)                                                                { eq.1 }

# (2) V1 + (3) V2 + (6) V3 = (100 V)

(2) {[(45) - (2) V2] / (2) } + (3) V2 + (6) V3 = (100 V)

(45) - (2) V2 + (3) V2 + (6) V3 = (100 V)

(1) V2 + (6) V3 = [(100 V) - (45 V)]

(1) V2 + (6) V3 = (55 V)                                                                     { eq.2 }

# (6.1) V2 + (7) V3 = (7.2 V)                                                              { eq.3 }

Equating eq.2 & 3, we get

(1) V2 + (6) V3 = (55 V)                                                            x 7

(6.1) V2 + (7) V3 = (7.2 V)                                                        x 6

-----------------------------------------------------------------------------------------------------------------------------------------

(7) V2 + (42) V3 = (385 V)

(36.6) V2 + (42) V3 = (43.2 V)

-----------------------------------------------------------------------------------------------------------------------------------------

- (29.6) V2 = (341.8 V)

V2 = - [(341.8 V) / (29.6)]

V2 = - 11.5 V

Inserting the value of 'V2' in eq.1, then we get

V1 = [(45 V) - (2) V2] / (2)  

V1 = [(45 V) - (2) (-11.5 V)] / (2)  

V1 = [(45 V) + (23 V)] / (2)  

V1 = [(68 V) / (2)]

V1 = 34 V